OpenStudy (shamil98):

clarification, Find the points on the curve where the tangent is horizontal or vertical. x = cos theta y = cos 3 theta so i know it's horizontal when dy/dtheta = 0 and i've found the points to be (1/2, -1) and (-1/2, 1) how would i determine when its vertical? dx/d theta = 0? the solutions manual says there is no vertical so i'm not sure..

OpenStudy (mathmale):

You're working with "parametric equations." If I remember correctly, the slope of a tangent line in such a situation is given by \[\frac{ dy }{ dx }=\frac{ \frac{ dy }{ d \theta } }{ \frac{ dx }{ d \theta } }\]

OpenStudy (mathmale):

where theta is the parameter in question. Please check this out in your textbook or online. This is not to say that your (dy / d theta) is wrong.

OpenStudy (mathmale):

Your tangent line would be vertical wherever the denominator (dx / d theta) is zero, since division by zero is undefined.

OpenStudy (mathmale):

I'd suggest that you find dy/dx as shown above in an equation involving the parameter theta. Reduce your result as much as possible. If it has a denominator, can that denom. every be zero?

OpenStudy (shamil98):

it would come to be the same thing would it not? dy/d theta = -3 sin 3 theta dx/d theta = -sin theta -3 sin 3 theta/- sin theta = 0 sin 3 theta = 0 at theta = pi/3, 2pi/3, -pi/3, -2pi/3 and for the vertical part dx/d theta = - sin theta theta = pi,2pi but my solutions manual says there is no vertical tangent

OpenStudy (mathmale):

I see y our point. Would it make a positive difference were you to expand sin (3 theta) into terms that include sin theta and then reduce the ensuing quotient?

OpenStudy (mathmale):

\[\sin 3 \theta = \sin (\theta + 2 \theta)=?\] Might work, might not, but perhaps worth a try.

OpenStudy (shamil98):

ah i hadn't thought of that, would make more sense if it were reduced that way and therefore there were no verticals at all..

OpenStudy (shamil98):

i came up with 3sin theta - 4sin^2 theta and yeah taking out a sin theta reduces the quotient to one :) my book went all the way on this one lol